OpenStudy (anonymous):

-6y+5x=10 gives me a 5/6 rise and run. how do i find the y intercept?

OpenStudy (anonymous):

replace x by 0 and solve

OpenStudy (anonymous):

get \[-6y+0=10\] \[y=-\frac{10}{6}=-\frac{5}{3}\]

OpenStudy (anonymous):

-6y+5x=10 subtract x by both sides -6y= -5x + 10 get y to be one by dividing -6y/-6y = y -5/-6 = 5/6 10/-6 = -10/6

OpenStudy (anonymous):

please break it down for me.

OpenStudy (anonymous):

i don't see how to gragh it on the calc.

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

well I dont know about graphing it on a calculator but i can break it down for you